# Prospects for the Detection of High-Redshift Gamma-Ray Bursts in the Era of EP and SVOM

### v1

@@ Abstract
Gamma-ray bursts (GRBs) are a promising probe of the high-redshift Universe, but their detection remains observationally challenging. In this work, we explore the detectability of high-$z$ GRBs by the Wide-field X-ray Telescope (WXT) aboard the Einstein Probe (EP) and the coded-mask gamma-ray imager (ECLAIRs) aboard the Space-based multi-band astronomical Variable Objects Monitor (SVOM). Using a population synthesis model calibrated to $Swift$ GRB observations, we develop a tool to estimate high-$z$ GRB detection rates for instruments with specific energy bands and sensitivities. Our results indicate that EP/WXT could detect $\sim 5.1^{+3.4}_{-2.4}$ (with 68% confidence level) GRBs annually at $z>6$, compared to $\sim 0.7^{+1.0}_{-0.4}$ $\mathrm{events\,yr^{-1}}$ at $z>6$ for SVOM/ECLAIRs. While EP cannot independently determine redshifts (requiring optical/near-infrared follow-up), its assumed $\sim 30\%$ follow-up efficiency yields $\sim 1.5^{+1.0}_{-0.7}$ confirmed $z>6$ GRBs annually. SVOM, equipped with dedicated follow-up telescopes, will ensure robust high-$z$ GRB confirmations. We anticipate that EP and SVOM will open new avenues for utilizing enlarged samples of high-$z$ GRBs to explore the early Universe. Moreover, EP will assemble a substantial sample of soft, low-luminosity GRBs at low-to-intermediate redshifts, providing critical insights into the structure of GRB jets.
@@ 1 Introduction
Gamma-ray bursts (GRBs), among the most energetic explosive transients in our Universe, are detectable up to very high redshifts. The most distant spectroscopically confirmed burst to date is GRB 090423 at $z=8.2$ (Salvaterra et al., 2009b; Tanvir et al., 2009). As such, GRBs represent a promising probe, complementary to quasar and galaxy surveys, for exploring the early Universe (see Salvaterra 2015; Wang et al. 2015 for reviews). In particular, long GRBs (defined by durations $T_{90}>2\,\mathrm{s}$) arise from the core collapse of massive stars (e.g., Woosley 1993; Paczyński 1998; Woosley & Bloom 2006), making them powerful tracers of star formation processes throughout cosmic history (e.g., Totani 1997; Chary et al. 2007; Yüksel et al. 2008; Kistler et al. 2009; Wang & Dai 2009; Wang 2013; Wei et al. 2014; Matsumoto et al. 2024). Moreover, GRB afterglows are expected to be observable up to $z\sim 20$ (Ciardi & Loeb, 2000; Gou et al., 2004), enabling spectroscopic afterglow studies to provide independent constraints on the reionization history, ionizing photon escape fraction, and cosmic metal enrichment history (Totani et al., 2006, 2014; Robertson & Ellis, 2012; Wang et al., 2012; Chornock et al., 2013; Hartoog et al., 2015; Fausey et al., 2025). GRBs may also constrain the stellar populations in the earliest galaxies by detecting explosions from the first, metal-free stars (Population III stars; Bromm & Loeb 2006; Mészáros & Rees 2010; de Souza et al. 2011; Suwa & Ioka 2011; Toma et al. 2011; Matsumoto et al. 2015, 2016; Kinugawa et al. 2019).
A major limitation is the extreme scarcity of high-redshift ($z\geq 6$) GRB detections, with only ten events having been identified in 20 years of Swift observations. Among these, seven have spectroscopic confirmations, while the remaining three rely on photometric redshift estimates (as listed in Table 1). This scarcity primarily results from current missions’ limited sensitivity and relatively high-energy bandpass, as high-$z$ GRBs are faint with peak emission redshifted to lower energies. Additionally, some high-$z$ GRBs may have been missed during the patchy follow-up campaigns. Such campaigns can be patchy because weather conditions sometimes prevented rapid follow-up observations, or afterglows were only observed at optical wavelengths and not in the near-infrared (Campana et al., 2022).
@@
Spectroscopy Photometry
@@
050904 6.3 1 090429B 9.4 12
080913 6.7 2 120521C 6.0 13
090423 8.2 3, 4 120923A 8.5 14
130606A 5.9 5, 6, 7
140515A 6.3 8, 9
210905A 6.3 10
240218A 6.8 11
References: (1) Kawai et al. (2006); (2) Greiner et al. (2009); (3) Salvaterra et al. (2009b); (4) Tanvir et al. (2009); (5) Chornock et al. (2013); (6) Totani et al. (2014); (7) Hartoog et al. (2015); (8) Chornock et al. (2014); (9) Melandri et al. (2015); (10) Saccardi et al. (2023); (11) Saccardi et al. (2024); (12) Cucchiara et al. (2011); (13) Laskar et al. (2014); (14) Tanvir (2013).
@@
In this work, we explore the prospects for high-$z$ GRB detection during the EP and SVOM observational era. EP was launched in January 2024, followed by SVOM in June 2024; both missions have now commenced scientific operations. Compared to other operational wide-field monitors, both the Wide-field X-ray Telescope (WXT) on board EP and the coded-mask gamma-ray imager (ECLAIRs) on board SVOM feature high sensitivity with coverage extending into relatively low X-ray energies. For a 10-s exposure, WXT achieves a sensitivity of $\sim 8.9\times 10^{-10}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ in 0.5–4 keV, while ECLAIRs reaches $\sim 7.2\times 10^{-8}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ in 4–150 keV. These properties make both instruments exceptionally well-suited for detecting high-$z$ GRBs. Notably, EP/WXT has already detected a GRB at $z=4.859$ (designated EP240315a; Liu et al. 2025), while SVOM/ECLAIRs achieved the fourth-most distant GRB redshift record with GRB 250314A at $z=7.3$ (Malesani et al., 2025), demonstrating both missions’ high-$z$ GRB survey capabilities.
@@ 2.2 Simulation Setup
The expected detection rate of GRBs for a detector with defined energy range and sensitivity becomes calculable when the intrinsic GRB formation rate, $\psi(z)$, and LF, $\phi(L)$, in Equation (3) are specified. However, these parameters remain poorly constrained. A key objective of GRB population studies is to infer them through statistical fitting of the observed distributions of redshift, luminosity, and peak flux measurements (e.g., Porciani & Madau 2001; Lloyd-Ronning et al. 2002, 2019; Firmani et al. 2004; Guetta et al. 2005; Natarajan et al. 2005; Daigne et al. 2006; Guetta & Piran 2007; Le & Dermer 2007; Salvaterra & Chincarini 2007; Kistler et al. 2008, 2009; Salvaterra et al. 2009a, 2012; Butler et al. 2010; Campisi et al. 2010; Qin et al. 2010; Wanderman & Piran 2010; Cao et al. 2011; Virgili et al. 2011; Lu et al. 2012; Robertson & Ellis 2012; Tan et al. 2013; Wei et al. 2014; Petrosian et al. 2015; Tan & Wang 2015; Yu et al. 2015; Deng et al. 2016; Pescalli et al. 2016; Wei & Wu 2017; Paul 2018; Lan et al. 2019, 2021; Le et al. 2020; Palmerio & Daigne 2021; Dong et al. 2022; Ghirlanda & Salvaterra 2022). Such analyses typically adopt parameterized models for both the GRB formation rate and LF, and iteratively refine them by matching simulated populations to observational data. Although the specific functional forms vary across studies, a consensus exists that GRB populations likely underwent some kind of redshift evolution (e.g., Salvaterra et al. 2009a, 2012; Virgili et al. 2011; Lan et al. 2019, 2021; Ghirlanda & Salvaterra 2022). Therefore, we employ several parameterized GRB rates and LFs, derived from recent representative population studies (Salvaterra et al., 2012; Lan et al., 2021; Ghirlanda & Salvaterra, 2022), to quantify the detection efficiency of high-$z$ GRBs for the EP/WXT and SVOM/ECLAIRs missions. Table 2 summarizes the energy bands, detection thresholds, fields of view, and duty cycles of these instruments.
@@
    where $\psi_{\star}(z)=(0.0157+0.118z)/[1+(z/3.23)^{4.66}]$ is the cosmoving cosmic star formation rate (SFR), with units of ${\rm M}_{\odot}$ $\rm yr^{-1}$ $\rm Mpc^{-3}$ (Hopkins & Beacom, 2006; Li, 2008). [^2] Note that a proportionality coefficient relating $\psi(z)$ (the GRB formation rate) to $\psi_{\star}(z)$ (the cosmic SFR) quantifies the GRB formation efficiency per unit stellar mass, with units of ${\rm M}^{-1}_{\odot}$. This coefficient is determined by normalizing the model predictions to the observed population of GRBs. S12 found that a strong redshift-dependent evolution in the GRB rate density ($\delta=1.7\pm 0.5$) is required to explain the observations.
@@ 3 Populations of High-z GRBs Accessible by EP and SVOM
Figure 3: Peak isotropic luminosity versus redshift of the simulated GRB populations detectable by EP/WXT (red solid contours) and SVOM/ECLAIRs (blue solid contours), with sensitivity thresholds for a 10-s exposure. The shaded regions represent the $1-3\sigma$ confidence contours. The blue dashed contours correspond to SVOM/ECLAIRs with a sensitivity threshold for a 1000-s exposure. The ten high-redshift ($z\geq 6$) GRBs detected by Swift are indicated by star symbols. A square symbol marks the detection of the long SVOM/ECLAIRs GRB 250314A at $z=7.3$.
By incorporating the properties of EP/WXT and SVOM/ECLAIRs (as listed in Table 2), we run our population synthesis code to perform Monte Carlo simulations. Figure 3 displays the contour levels in the plane of the peak isotropic luminosity $L$ versus redshift $z$ for the simulated GRB populations detectable by EP/WXT (red solid contours) and SVOM/ECLAIRs (blue solid contours). For comparison, we include the ten high-redshift ($z\geq 6$) GRBs detected by Swift (star symbols). Due to its extended energy range into soft X-rays and higher sensitivity, the EP/WXT simulations sample relatively less energetic events compared to those currently detected by Swift and those expected from SVOM in the near future. The detection of the long SVOM/ECLAIRs GRB 250314A at $z=7.3$ (square symbol; Malesani et al. 2025) appears to lie outside the $3\sigma$ confidence region predicted by SVOM/ECLAIRs simulations (blue solid contours). In our simulations, the fiducial sensitivity of SVOM/ECLAIRs is set to $7.2\times 10^{-8}$ $\mathrm{erg\,cm^{-2}\,s^{-1}}$, achievable with an exposure time of $\simeq 10\,\mathrm{s}$. The instrument’s maximal sensitivity, however, reaches $7.2\times 10^{-10}$ $\mathrm{\rm erg\,cm^{-2}\,s^{-1}}$ for exposures of $\simeq 1000\,\mathrm{s}$ (Wei et al., 2016). Adopting this higher sensitivity threshold, the simulated GRB population detectable by SVOM/ECLAIRs shifts to the blue dashed contours. Notably, GRB 250314A now falls within the $3\sigma$ confidence region of these updated predictions.
@@ 4 Summary and Discussions
Note that EP, operating solely as an X-ray telescope, cannot determine redshifts. Optical/near-infrared follow-up observations are essential for identifying high-$z$ GRBs, though such follow-ups may not always be feasible for EP-detected events. This could yield fewer confirmed high-$z$ GRBs than our estimates suggest. Assuming a $\sim 30\%$ efficiency for ground-based spectroscopic redshift determination, we expect EP to detect $\sim 1.5^{+1.0}_{-0.7}$ GRBs at $z>6$ annually. Equipped with dedicated follow-up telescopes, SVOM will play a crucial role in high-$z$ GRB detection and confirmation (see Llamas Lanza et al. 2024 for details).
After several years of operation, we anticipate that EP and SVOM will double the number of existing high-$z$ GRBs. Furthermore, their substantial sample of low-luminosity GRBs at low-to-intermediate redshifts will enable investigations into potential sub-populations with dominant soft X-ray flashes and the faint end of the GRB LF—both offering key insights into GRB jet structure (Liang et al., 2007; Pescalli et al., 2015; Salafia et al., 2016). In conclusions, EP and SVOM will open new avenues for advancing GRB studies and early Universe exploration through high-$z$ beacons that probe the cosmic frontier.

### v2

@@ Abstract
Gamma-ray bursts (GRBs) are a promising probe of the high-redshift Universe, but their detection remains observationally challenging. In this work, we explore the detectability of high-$z$ GRBs by the Wide-field X-ray Telescope (WXT) aboard the Einstein Probe (EP) and the coded-mask gamma-ray imager (ECLAIRs) aboard the Space-based multi-band astronomical Variable Objects Monitor (SVOM). Using a population synthesis model calibrated to $Swift$ GRB observations, we develop a tool to estimate high-$z$ GRB detection rates for instruments with specific energy bands and sensitivities. Our results indicate that EP/WXT could detect $\sim 5.1^{+3.4}_{-2.4}$ (with 68% confidence level) GRBs annually at $z>6$, compared to $\sim 0.7^{+1.0}_{-0.4}$ $\mathrm{events\,yr^{-1}}$ at $z>6$ for SVOM/ECLAIRs. While EP cannot independently determine redshifts (requiring optical/near-infrared follow-up), its assumed $\sim 30\%$ follow-up efficiency yields $\sim 1.5^{+1.0}_{-0.7}$ confirmed $z>6$ GRBs annually. SVOM, equipped with dedicated follow-up telescopes, will promptly identify high-$z$ candidates deserving deep near-infrared spectroscopy to ensure robust confirmation of high-$z$ GRBs. We anticipate that EP and SVOM will open new avenues for utilizing enlarged samples of high-$z$ GRBs to explore the early Universe. Moreover, EP will assemble a substantial sample of soft, low-luminosity GRBs at low-to-intermediate redshifts, providing critical insights into the structure of GRB jets.
@@ 1 Introduction
Gamma-ray bursts (GRBs), among the most energetic explosive transients in our Universe, are detectable up to very high redshifts. The most distant spectroscopically confirmed burst to date is GRB 090423 at $z=8.2$ (Salvaterra et al., 2009b; Tanvir et al., 2009). As such, GRBs represent a promising probe, complementary to quasar and galaxy surveys, for exploring the early Universe (see Salvaterra 2015; Wang et al. 2015 for reviews). In particular, long GRBs with durations $T_{90}>2\,\mathrm{s}$ (where $T_{90}$ is defined as the time interval containing 90% of the prompt emission; Kouveliotou et al. 1993) arise from the core collapse of massive stars [^1] Notably, two nearby long-duration bursts, GRB 211211A and GRB 230307A, exhibit kilonova signatures indicative of compact binary mergers, implying that at least some long GRBs originate from non-collapsar progenitors (Rastinejad et al., 2022; Troja et al., 2022; Yang et al., 2022; Dichiara et al., 2023; Levan et al., 2024). Given that merging compact binary progenitors have significantly longer average lifetimes than collapsars, the fraction of merger-origin long-duration GRBs (especially those bright enough to be detectable) is expected to decrease substantially at high redshifts compared to $z=0$. (e.g., Woosley 1993; Paczyński 1998; Woosley & Bloom 2006), establishing them as powerful tracers of cosmic star formation (e.g., Totani 1997; Chary et al. 2007; Yüksel et al. 2008; Kistler et al. 2009; Wang & Dai 2009; Robertson & Ellis 2012; Wang 2013; Wei et al. 2014; Matsumoto et al. 2024). Moreover, GRB afterglows are expected to be observable up to $z\sim 20$ (Ciardi & Loeb, 2000; Gou et al., 2004), enabling spectroscopic afterglow studies to provide independent constraints on the reionization history, ionizing photon escape fraction, and cosmic metal enrichment history, and $\mathrm{21\,cm}$ absorption features (Ioka & Mészáros, 2005; Totani et al., 2006, 2014; Wang et al., 2012; Chornock et al., 2013, 2014; Ciardi et al., 2015; Hartoog et al., 2015; Tanvir et al., 2019; Saccardi et al., 2023; Fausey et al., 2025a, b; Šoltinský et al., 2025). GRBs may also constrain the stellar populations in the earliest galaxies by detecting explosions from the first, metal-free stars (Population III stars; Bromm & Loeb 2006; Mészáros & Rees 2010; de Souza et al. 2011; Suwa & Ioka 2011; Toma et al. 2011; Matsumoto et al. 2015, 2016; Kinugawa et al. 2019).
A major limitation is the extreme scarcity of high-redshift ($z\geq 6$) GRB detections, with only eleven events having been identified in 20 years of Swift observations. Among these, seven have spectroscopic confirmations, while the remaining four rely on photometric redshift estimates (as listed in Table 1). This scarcity primarily results from current missions’ limited sensitivity and relatively high-energy bandpass, as high-$z$ GRBs are faint with peak emission redshifted to lower energies. Although Swift is suboptimal for detecting high-$z$ GRBs, previous studies suggest that $\sim 2\%$ of its GRBs originate at $z>5.5$ (Perley et al., 2016). However, some high-$z$ events may have been missed during the patchy follow-up campaigns. Such patchiness arises when weather conditions prevent rapid follow-up observations, when afterglows are only observed at optical wavelengths rather than in the near-infrared (NIR), and because ground facilities with deep NIR spectroscopic capabilities for high-$z$ afterglows (requiring 8 m-class telescopes) remain scarce (Campana et al., 2022).
@@
Spectroscopic Redshifts Photometric Redshifts
@@
050904 6.3 (1) 090429B 9.4 (12)
080913 6.7 (2) 100905A 7.9 (13)
090423 8.2 (3), (4) 120521C 6.0 (14)
130606A 5.9 (5), (6), (7) 120923A 7.8 (15)
140515A 6.3 (8), (9)
210905A 6.3 (10)
240218A 6.8 (11)
References: (1) Kawai et al. (2006); (2) Greiner et al. (2009); (3) Salvaterra et al. (2009b); (4) Tanvir et al. (2009); (5) Chornock et al. (2013); (6) Totani et al. (2014); (7) Hartoog et al. (2015); (8) Chornock et al. (2014); (9) Melandri et al. (2015); (10) Saccardi et al. (2023); (11) Saccardi et al. (2024); (12) Cucchiara et al. (2011); (13) Bolmer et al. (2018); (14) Laskar et al. (2014); (15) Tanvir et al. (2018).
@@
In this work, we explore the prospects for high-$z$ GRB detection during the EP and SVOM observational era. EP was launched in January 2024, followed by SVOM in June 2024; both missions have now commenced scientific operations. Compared to other operational wide-field monitors, both the Wide-field X-ray Telescope (WXT) on board EP and the coded-mask gamma-ray imager (ECLAIRs) on board SVOM feature high sensitivity with coverage extending into relatively low X-ray energies. For a 10-s exposure, WXT achieves a sensitivity of $\sim 8.9\times 10^{-10}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ in 0.5–4 keV, while ECLAIRs reaches $\sim 7.2\times 10^{-8}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ in 4–150 keV. These properties make both instruments exceptionally well-suited for detecting high-$z$ GRBs. Notably, EP/WXT has already detected a GRB at $z=4.859$ (designated EP240315a; Liu et al. 2025), while SVOM/ECLAIRs achieved the fifth-most distant GRB redshift record with GRB 250314A at $z=7.3$ (Malesani et al., 2025), demonstrating both missions’ high-$z$ GRB survey capabilities.
@@ 2.2 Simulation Setup
The expected detection rate of GRBs for a detector with defined energy range and sensitivity becomes calculable when the intrinsic GRB formation rate, $\psi(z)$, and LF, $\phi(L)$, in Equation (3) are specified. However, these parameters remain poorly constrained. A key objective of GRB population studies is to infer them through statistical fitting of the observed distributions of redshift, luminosity, and peak flux measurements (e.g., Porciani & Madau 2001; Lloyd-Ronning et al. 2002, 2019; Firmani et al. 2004; Guetta et al. 2005; Natarajan et al. 2005; Daigne et al. 2006; Guetta & Piran 2007; Le & Dermer 2007; Salvaterra & Chincarini 2007; Kistler et al. 2008, 2009; Salvaterra et al. 2009a, 2012; Butler et al. 2010; Campisi et al. 2010; Qin et al. 2010; Wanderman & Piran 2010; Cao et al. 2011; Virgili et al. 2011; Lu et al. 2012; Robertson & Ellis 2012; Tan et al. 2013; Wei et al. 2014; Petrosian et al. 2015; Tan & Wang 2015; Yu et al. 2015; Deng et al. 2016; Pescalli et al. 2016; Wei & Wu 2017; Paul 2018; Lan et al. 2019, 2021; Le et al. 2020; Palmerio & Daigne 2021; Dong et al. 2022; Ghirlanda & Salvaterra 2022; Qu et al. 2024, 2025). Such analyses typically adopt parameterized models for both the GRB formation rate and LF, and iteratively refine them by matching simulated populations to observational data. Although the specific functional forms vary across studies, a consensus exists that GRB populations likely underwent some kind of redshift evolution (e.g., Salvaterra et al. 2009a, 2012; Virgili et al. 2011; Lan et al. 2019, 2021; Ghirlanda & Salvaterra 2022). Therefore, we employ several parameterized GRB rates and LFs, derived from recent representative population studies (Salvaterra et al., 2012; Lan et al., 2021; Ghirlanda & Salvaterra, 2022), to quantify the detection efficiency of high-$z$ GRBs for the EP/WXT and SVOM/ECLAIRs missions. Table 2 summarizes the energy bands, detection thresholds, fields of view, and duty cycles of these instruments.
@@
    where $\psi_{\star}(z)=(0.0157+0.118z)/[1+(z/3.23)^{4.66}]$ is the comoving cosmic star formation rate (SFR), with units of ${\rm M}_{\odot}$ $\rm yr^{-1}$ $\rm Mpc^{-3}$ (Hopkins & Beacom, 2006; Li, 2008). [^3] Note that a proportionality coefficient relating $\psi(z)$ (the GRB formation rate) to $\psi_{\star}(z)$ (the cosmic SFR) quantifies the GRB formation efficiency per unit stellar mass, with units of ${\rm M}^{-1}_{\odot}$. This coefficient is determined by normalizing the model predictions to the observed population of GRBs. S12 found that a strong redshift-dependent evolution in the GRB rate density ($\delta=1.7\pm 0.5$) is required to explain the observations.
@@ 3 Populations of High-z GRBs Accessible by EP and SVOM
Figure 3: Peak isotropic luminosity versus redshift of the simulated GRB populations detectable by EP/WXT (red solid contours) and SVOM/ECLAIRs (blue solid contours), with sensitivity thresholds for a 10-s exposure. The shaded regions represent the $1-3\sigma$ confidence contours. The blue dashed contours correspond to SVOM/ECLAIRs with a sensitivity threshold for a 1000-s exposure. The eleven high-redshift ($z\geq 6$) GRBs detected by Swift are indicated by star symbols. A square symbol marks the detection of long GRB 250314A at $z=7.3$ by SVOM/ECLAIRs.
By incorporating the properties of EP/WXT and SVOM/ECLAIRs (as listed in Table 2), we run our population synthesis code to perform Monte Carlo simulations. Figure 3 displays the contour levels in the plane of the peak isotropic luminosity $L$ versus redshift $z$ for the simulated GRB populations detectable by EP/WXT (red solid contours) and SVOM/ECLAIRs (blue solid contours). For comparison, we include the eleven high-redshift ($z\geq 6$) GRBs detected by Swift (star symbols). [^7] The relevant information for the peak luminosities of these eleven Swift GRBs is available from published literature. For GRB 210905A, we used $L=1.87\times 10^{53}$ $\mathrm{erg\,s^{-1}}$ from Rossi et al. (2022). For GRB 240218A, we calculated $L=2.88\times 10^{53}$ $\mathrm{erg\,s^{-1}}$ using the peak photon flux and spectral parameters reported in Veres et al. (2024). For the other nine GRBs, peak luminosities were taken from Lan et al. (2021). Due to its extended energy range into soft X-rays and higher sensitivity, the EP/WXT simulations sample relatively less energetic events compared to those currently detected by Swift and those expected from SVOM in the near future. GRB 250314A at $z=7.3$, [^8] The peak luminosity of GRB 250314A is $L=1.97\times 10^{52}$ $\mathrm{erg\,s^{-1}}$, based on private communication with the SVOM team. detected by SVOM/ECLAIRs (square symbol; Malesani et al. 2025), lies outside the $3\sigma$ confidence region predicted by SVOM/ECLAIRs simulations (blue solid contours). In our simulations, the fiducial sensitivity of SVOM/ECLAIRs is set to $7.2\times 10^{-8}$ $\mathrm{erg\,cm^{-2}\,s^{-1}}$, achievable with an exposure time of $\simeq 10\,\mathrm{s}$. The instrument’s maximal sensitivity, however, reaches $7.2\times 10^{-10}$ $\mathrm{\rm erg\,cm^{-2}\,s^{-1}}$ for exposures of $\simeq 1000\,\mathrm{s}$ (Wei et al., 2016). Adopting this higher sensitivity threshold, the simulated GRB population detectable by SVOM/ECLAIRs shifts to the blue dashed contours. Notably, GRB 250314A now falls within the $3\sigma$ confidence region of these updated predictions. In the 15–150 keV energy band, Swift-detected long GRBs typically last tens to hundreds of seconds (Horváth & Tóth, 2016). At lower energies ($\mathrm{<10\,keV}$), they show substantially longer durations, with EP-observed GRBs averaging several hundred seconds (Gao et al., 2025). While cosmological time-dilation would prolong observed durations at high redshifts, bandpass-shifting and undetected weak prompt emission episodes counteract this effect (Littlejohns et al., 2013). These factors collectively justify longer exposures for soft X/gamma-ray instruments like EP/WXT and SVOM/ECLAIRs.
@@ 4 Summary and Discussions
Note that EP, operating solely as an X-ray telescope, cannot determine redshifts. Optical/NIR follow-up observations are essential for identifying high-$z$ GRBs, though such follow-ups may not always be feasible for EP-detected events. This could yield fewer confirmed high-$z$ GRBs than our estimates suggest. Assuming a $\sim 30\%$ efficiency for ground-based spectroscopic redshift determination, we expect EP to detect $\sim 1.5^{+1.0}_{-0.7}$ GRBs at $z>6$ annually. Equipped with dedicated follow-up telescopes, SVOM will play a crucial role in promptly identifying high-$z$ GRB candidates deserving deep NIR spectroscopy (see Llamas Lanza et al. 2024). Specifically, very high-redshift cases may be indicated by non-detection in both channels of the on-board Visible Telescope (VT) (though not uniquely, as dusty GRBs would also appear optically faint). These candidates can then be prioritized for ground-based NIR follow-up observations using positional locations from SVOM’s Microchannel X-ray Telescope (MXT).
After several years of operation, we anticipate that EP and SVOM will double the number of existing high-$z$ GRBs. Furthermore, their substantial sample of low-luminosity GRBs at low-to-intermediate redshifts will enable investigations into potential sub-populations with dominant soft X-ray flashes and the faint end of the GRB LF—both offering key insights into GRB jet structure (Liang et al., 2007; Pescalli et al., 2015; Salafia et al., 2016). In conclusions, EP and SVOM will open new avenues for advancing GRB studies and early Universe exploration through high-$z$ beacons that probe the cosmic frontier. Notably, besides EP and SVOM, several mission concepts specifically designed for high-$z$ GRB discovery, including THESEUS (Amati et al., 2018, 2021; Ghirlanda et al., 2021), Gamow Explorer (White, 2020), and HiZ-GUNDAM (Yonetoku et al., 2024), are under active investigation.